6-)Find f(x)=(x-2)^3(x-4),
accelerate newton iteration numerically
sand show that R=2
Answer #1

For the given function
f(x) = (x - 2)^3 *(x - 4)
so,
f'(x) = 3(x - 2)^2(x - 4) + (x - 2)^3

now to find roots we have to put f(x) = 0
from there we can easily see x = 2 and x = 4 are solutions and R = 2 is repeated root

now, from newtons method, let xo = 1
then x1 = xo - f(xo)/f'(xo)
x1 = 1 - (-1)(-3)/(3(-3) + 1) = 1 + 3/8 = 11/8
the following excel sheet has values till x18

x0 1.00000
x1 1.30000
x2 1.51477
x3 1.66663
x4 1.77270
x5 1.84598
x6 1.89612
x7 1.93019
x8 1.95320
x9 1.96868
x10 1.97907
x11 1.98602
x12 1.99067
x13 1.99378
x14 1.99585
x15 1.99723
x16 1.99815
x17 1.99877
x18

1.99918

taking initial value 3 gives us

x0 3.00000
x1 2.50000
x2 2.31250
x3 2.20148
x4 2.13172
x5 2.08675
x6 2.05739
x7 2.03807
x8 2.02530
x9 2.01683
x10 2.01120
x11 2.00746
x12 2.00497
x13 2.00331
x14 2.00221
x15 2.00147
x16 2.00098
x17 2.00065
x18 2.00044

using initial value 3.5 gives us division by 0 error, using 3.6 gives

x0 3.60000
x1 5.20000
x2 4.63529
x3 4.26663
x4 4.06955
x5 4.00637
x6 4.00006
x7 4.00000
x8 4.00000
x9 4.00000
x10 4.00000
x11 4.00000
x12 4.00000
x13 4.00000
x14 4.00000
x15 4.00000
x16 4.00000
x17 4.00000
x18 4.00000

hence we can see, R = 2, is repeated root and E = 4 is single root